# Scalarisation-based risk concepts for robust multi-objective optimisation

### v2

@@ 4.2 Risk statistics
with $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\xi})]=\sup_{\mathbf{x}\in\mathbb{X}}s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(f(\mathbf{x},\boldsymbol{\xi}))$, for any uncertain vector $\boldsymbol{\xi}\in\Xi$ and positive unit vector $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. If the objective function $f$ is bounded, then this set is either equal to the Pareto front surface $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\{f(\mathbf{x},\boldsymbol{\xi})\}_{\mathbf{x}\in\mathbb{X}}]$, if it is non-empty, or it is the degenerate singleton set $\{\boldsymbol{\eta}\}\in\mathbb{L}_{\boldsymbol{\eta}}$. As shown by Tu et al. (2024b, Section 4), one can readily use this polar representation in order to generalise standard statistics from the univariate setting to function over the space of Pareto front surfaces. For example, one can define the expected Pareto front surface as the polar surface constructed using the expected projected lengths (Example 4.1).
@@
As shown by Tu et al. (2024b, Section 4.1), if the objective function $f$ is bounded and the Pareto front surfaces $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\{f(\mathbf{x},\boldsymbol{\xi})\}_{\mathbf{x}\in\mathbb{X}}]$ are non-empty almost surely, then this polar surface is a valid Pareto front surface. Otherwise it could degenerate to the singleton set $\{\boldsymbol{\eta}\}\in\mathbb{L}_{\boldsymbol{\eta}}$.
We now present a straightforward extension of this result in Proposition 4.1 which holds for any partially coherent univariate risk functional. The proof of this result is presented in Section A.6.
@@
for some positive directions $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$, where $P>0$ denotes a cardinality constraint. Individually, each of these problems tries to identify a collection of inputs $X\subseteq\mathbb{X}$, whose collective output, under the preference direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$, is of high quality. This problem reduces down to the STR length scalarised problem (12) when $P=1$. Notably, this is a monotone and submodular set optimisation problem when $P>1$ and therefore is, in general, much more computationally challenging to solve than the corresponding RTS or STR problems—see Tu et al. (2024a) for more details.
@@ 4.3 Robust Pareto front surfaces
In Proposition 4.2 and Proposition 4.3, we present the main results in this section which states that these polar surfaces (20) and (21) are indeed valid Pareto front surfaces. Also, we give an explicit representation on what these Pareto front surfaces actually are. In words, the RTS front is the Pareto front surface associated with the union robust problem (10). Whilst the STR front is the Pareto front surface associated with the risk statistic of each input’s Pareto front surface distribution. The proof of these result are presented in Section A.7 and Section A.8, respectively.
@@
In Figure 5, we illustrate these risk-adjusted Pareto front surfaces for a simple two-dimensional example with the risk functionals introduced in Section 2.3. We also compare these Pareto front surfaces with the risk statistics described in Proposition 4.1. Overall, we see a great variety between the Pareto front surfaces illustrated within each plot; the differences among the plots however are very subtle. Conceptually, the Pareto front surface statistics looks at the aggregated performance at each scalarised value and therefore has a noticeably smoother surface. In contrast, the RTS and STR front looks at the individual performance at each scalarised level and therefore is a bit more bumpy. In particular, the RTS front treats every input $\mathbf{x}\in X$ according to the Pareto front surface of its risk-adjusted output set $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\boldsymbol{\rho}[f(\mathbf{x},\cdot)]]$. Whilst the STR front treats every input $\mathbf{x}\in X$ according to a risk statistic of its random Pareto front surface $\rho[\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[f(\mathbf{x},\cdot)]]$. These two formulations are different in general. However, as highlighted in Section 3.3.3, there are special cases when they are equivalent, such as: the worst-case setting (blue lines) and the value-at-risk setting (orange lines).
Figure 5: An illustration of the different Pareto front surfaces that can be computed from a two-dimensional objective function. On the left, we plot the output sets $f(\mathbf{x},\Xi)\subset\mathbb{R}^{M}$ for each input $\mathbf{x}\in\mathbb{X}$, which are all assumed to be uniformly distributed in some ellipsoid. On the remaining plots, we draw the Pareto front surfaces associated with the Pareto front statistics, RTS fronts and STR fronts for the different risk functionals described in Section 2.3.
@@ 5 Robust performance metrics
Performance metrics are often used in multi-objective optimisation in order to quantify the quality of a Pareto front approximation. As showcased in a recent survey by Audet et al. (2021), there exists many different types of performance metrics in the multi-objective literature. All of which try to measure different facets of the Pareto front approximation. As established by Tu et al. (2024a), the most natural class of performance metrics, under the scalarisation perspective, are the family of R2 utilities originally proposed by Hansen and Jaszkiewicz (1998). Notably, this family of performance metrics exhibit many favourable properties and contains many of the most widely-used performance metrics as special cases (Tu et al., 2024a). We will now generalise this family of utilities to the robust setting via the RTS and STR methodology. In particular, we will illustrate some robust generalisations to the well-known inverted generational distance (Section 5.2) and the hypervolume indicator (Section 5.3).
@@ 5.1 Robust R2 utilities
Formally, a multi-objective performance metric is any utility function $U:2^{\mathbb{R}^{M}}\rightarrow\mathbb{R}$, which maps a set of vectors to a scalar. We adopt the convention that larger utility values signify better performance. The R2 utilities is a special class of utility functions that is defined with respect to a family of scalarisation functions $\{s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}\}_{\boldsymbol{\theta}\in\Theta}$ and a probability distribution over the corresponding scalarisation parameters $p\in\mathcal{P}[\Theta]$. Specifically, equipped with this set-up, the corresponding R2 utility of a set of vectors $Y\subset\mathbb{R}^{M}$ is equal to the weighted average maximum scalarised value
@@
Note that this is the general form of an R2 utility which also accounts for the setting where the Pareto front approximation $Y\subset\mathbb{R}^{M}$ is continuous. Nevertheless, in practice, the Pareto front approximation is typically finite and therefore the supremum operation is often replaced with the maximum. At a intuitive level, an R2 utility just computes a weighted score associated with a collection of scalarised problems (3). By design, this family of utility functions is very flexible and it gives us any easy way to define and measure different notions of performance in the objective space. In theory, one could take advantage of this flexibility in order to define R2 utilities which adequately cater for the decision maker’s internal preferences.
@@
Inspired by the construction of the R2 utilities (22), we now define the corresponding RTS and STR extension of the R2 utilities by replacing the maximum scalarised values arising in (3) with the RTS scalarised values (8) and STR scalarised values (11), respectively. Precisely, we define an RTS-R2 utility function $R_{f,\boldsymbol{\rho}}^{\text{RTS}}:2^{\mathbb{X}}\rightarrow\mathbb{R}$ and a STR-R2 utility function $R_{f,\rho}^{\text{STR}}:2^{\mathbb{X}}\rightarrow\mathbb{R}$ by the following equations:
@@
Notice that one can rewrite an RTS-R2 utility in terms of a standard R2 utility: $R^{\textnormal{RTS}}_{f,\boldsymbol{\rho}}[X]=U[\cup_{\mathbf{x}\in X}\boldsymbol{\rho}[f(\mathbf{x},\cdot)]]$. Therefore, to compute the RTS performance of a set of inputs, one can just has to compute the corresponding R2 utility of the union risk-adjusted output set. In general, a similar result does not hold for the STR-R2 utility unless the operations commute (Section 3.3.3).
@@ 5.2 Robust inverted generational distance
The inverted generational distance (IGD) is a popular multi-objective performance metric that was initially introduced by Coello and Reyes Sierra (2004). Following the Schutze et al. (2012), the IGD indicator $I^{\text{IGD}_{p,q}}:2^{\mathbb{R}^{M}}\times B[\mathbb{R}^{M}]\rightarrow\mathbb{R}$ computes some notion of the average distance between a set of vectors $Y\subset\mathbb{R}^{M}$ and a finite set of ideal vectors $\Upsilon\in\mathbb{B}[\mathbb{R}^{M}]$,
@@
for some norms $p,q\geq 1$, where $\mathbb{B}[\mathbb{R}^{M}]:=\{Y\subset\mathbb{R}^{M}:|Y|<\infty\}$ denotes the space of finite sets of vectors. By construction, a smaller IGD indicator value is more desirable. Following Tu et al. (2024a), we can rewrite the IGD indicator as a transformation of an R2 utility, namely the IGD utility
@@
which holds for any set of inputs $X\subseteq\mathbb{X}$ and norms $p,q\geq 1$. In words, the RTS-IGD utility computes the average negative distance between the risk-adjusted output set and the set of utopia points. Whilst, the STR-IGD utility computes the average risk-adjusted negative distance between the distribution of possible objective vectors and the set of utopia points. On the left of Figure 6, we give a pictorial illustration of these different computations on a simple two-dimensional example.
@@
We can also generalise the above discussion for the corresponding IGD+ indicator described by Ishibuchi et al. (2015). This is a straightforward extension, where the IGD scalarisation function is replaced with its truncated version (Tu et al., 2024a).
@@ 5.3 Robust hypervolume indicator
for any set of inputs $X\subseteq\mathbb{X}$, bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, partially coherent univariate risk functional $\rho$ and reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. By a simple calculation, we see that the variance of this estimate can uniformly upper bounded by a function dependent constant divided by the number of Monte Carlo samples, that is
@@ 6 Numerical examples
We now present two numerical case studies in order to highlight the efficacy of the novel ideas that we have presented in this work. Firstly, in Section 6.1, we present the cake baking case study which illustrates a practical example where one has to decide between an RTS and STR solution. Whilst secondly, in Section 6.2, we present the robust rocket injector case study which showcases how one can use the concepts of a robust Pareto front and a robust performance metric in order to devise and assess the performance of some robust multi-objective optimisation algorithms.
@@ 6.1 Cake problem
changes when the relative composition of its $D=6$ main ingredients are varied. The inputs for this problem are the relative amount of ingredients $\mathbf{x}\in\mathbb{X}$, which lie in the $(D-1)$-dimensional simplex and satisfies the following inequality constraints:
@@
Suppose for instance that the decision maker is interested in identifying the ingredients that is required to make a cake whose properties achieve a target value of $\boldsymbol{\upsilon}=(-360,4,1)\in\mathbb{R}^{M}$. As motivated earlier in Example 3.1, one can recast this target optimisation problem as a robust scalarised optimisation problem. Specifically, one can use the expectation functionals (Examples 2.3 and 2.8) along with a distance-based scalarisation function. As a concrete example, we consider using the following weighted variant of the IGD scalarisation function:
@@ 6.2 Robust rocket injector problem
For the manufacturer’s distribution, we assume a uniform distribution $p(\boldsymbol{\xi})\sim\text{Uniform}(\Xi)$ over the uncertain parameters $\boldsymbol{\xi}\in\Xi$. Whilst, for our robust objective, we consider an STR approach where the goal is to identify the STR robust designs whose output distribution lies on the CVaR Pareto front surface at level $\alpha=0.9$. Equivalently, we formulate this problem as a robust utility optimisation problem where we are interested in optimising the STR hypervolume (Proposition 5.2) of the identified CVaR front
@@
To emulate a real-world scenario, we treat the objective function $f$ as a black-box function that can only be evaluated sparingly at a few carefully selected design locations $T=65$. Precisely, at any time $t>0$, the decision maker is given the opportunity to see a noisy function evaluation at any chosen design $\mathbf{x}_{t}\in\mathbb{X}$, that is $\mathbf{y}_{t}=f(\mathbf{x}_{t},\boldsymbol{\xi}_{t})+\boldsymbol{\epsilon}_{t}\in\mathbb{R}^{M}$, where $\boldsymbol{\xi}_{t}\sim p(\boldsymbol{\xi})$ is the uncertain parameter that was independently sampled by the manufacturer and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{M},\textsc{Diag}(\boldsymbol{\sigma}^{2}))$ is the observation noise. In our example, we consider a zero-mean Gaussian observation noise whose standard deviation $\boldsymbol{\sigma}=(0.01,0.01,0.01)$ is set to $1\%$ of the estimated scaled objective range. To initialise this problem, we randomly sampled $T_{0}=5$ initial points from the controllable input space. In Figure 9, we illustrate these initial points in the objective space and plot the corresponding CVaR front under the STR approach. Notably, to compute this CVaR front, we require knowledge of the true objective function at each uncertain parameter which is typically not available in an actual real-world experimental set-up.
@@
where $X_{t-1}=\{\mathbf{x}_{n}\}_{n\leq t-1}$ denotes the set of previously queried inputs, whilst $\hat{\boldsymbol{\eta}}_{t}\in\mathbb{R}^{M}$ and $\hat{f}_{t-1}:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$ denotes the latest estimate of the reference vector and objective function, respectively. In our example, we set the latest reference vector to be an over-estimation of the lower bound $\eta_{t-1}^{(m)}=\min_{n<t}y_{n}^{(m)}-\kappa(\max_{n<t}y_{n}^{(m)}-\min_{n<t}y_{n}^{(m)})$ with $\kappa=0.1$ for $m=1,\dots,M$. Whilst for the surrogate model, we used an upper confidence bound of a Gaussian process model which was trained on the previous observations, that is $\hat{f}^{\text{UCB}}_{t-1}(\mathbf{x},\boldsymbol{\xi})=\boldsymbol{\mu}_{t-1}(\mathbf{x},\boldsymbol{\xi})+\beta\boldsymbol{\sigma}_{t-1}(\mathbf{x},\boldsymbol{\xi})$, where $\beta=2.0$ is the trade-off parameter and the vectors $\boldsymbol{\mu}_{t-1}(\mathbf{x},\boldsymbol{\xi})\in\mathbb{R}^{M}$ and $\boldsymbol{\sigma}_{t-1}(\mathbf{x},\boldsymbol{\xi})\in\mathbb{R}_{\geq 0}^{M}$ are the mean and standard deviation of the model at the input pair $(\mathbf{x},\boldsymbol{\xi})\in\mathbb{X}\times\Xi$. We refer to this overall acquisition strategy as the STR-HVI-UCB approach.
@@
Figure 10: A numerical comparison between the different multi-objective algorithms on the robust rocket injector problem over $100$ independent runs. For each method, we plot the mean log regret at every iteration along with an uncertainty estimate. Precisely, we highlight the band between the $0.25$ to $0.75$ and $0.1$ to $0.9$ quantiles with a dark and light shade, respectively.
In Figure 10, we present the results of our numerical comparison in which we compared our proposed method versus the existing baselines over $100$ independent runs on a discretised version of the robust rocket injector problem. Namely we compare our method against non-robust approaches such as: the random search (Sobol), NSGA2 (Deb et al., 2002), NParEGO (Knowles, 2006; Daulton et al., 2021), NEHVI (Emmerich et al., 2006; Daulton et al., 2021), PFES (Suzuki et al., 2020) and the JES-LB (Tu et al., 2022) algorithms. We also compare our approach against the MARS algorithm (Daulton et al., 2022), which is a robust algorithm that takes an STR approach using the Chebyshev scalarisation function and the univariate value-at-risk risk functional. Unsurprisingly we see that the robust methods perform better than the non-robust methods. Moreover, our method performs the best overall under the chosen utility. For some visual intuition, we have included Figure 11 where we plot the final CVaR front obtained by each algorithm under the same random seed. Visually we see that most of the variation between these algorithms transpire around the first and second objectives. The existing approaches regularly struggle to find the inputs which robustly maximises the CVaR front over these two objectives. In contrast, our robust approach is much more consistent and is often very quick in identifying the difficult inputs which robustly maximises the whole CVaR front.

### v4

@@ 4.2 Risk statistics
with $\ell_{\boldsymbol{\eta},\boldsymbol{\lambda}}[Y_{\boldsymbol{\eta},f}^{*}(\boldsymbol{\xi})]=\sup_{\mathbf{x}\in\mathbb{X}}s^{\textnormal{Len}}_{(\boldsymbol{\eta},\boldsymbol{\lambda})}(f(\mathbf{x},\boldsymbol{\xi}))$, for any uncertain vector $\boldsymbol{\xi}\in\Xi$ and positive unit vector $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$. If the objective function $f$ is bounded, then this set is either equal to the Pareto front surface $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\{f(\mathbf{x},\boldsymbol{\xi})\}_{\mathbf{x}\in\mathbb{X}}]$, if it is non-empty, or it is the degenerate singleton set $\{\boldsymbol{\eta}\}\in\mathbb{L}_{\boldsymbol{\eta}}$. As shown by Tu et al. (2024, Section 4), one can readily use this polar representation in order to generalise standard statistics from the univariate setting to the space of Pareto front surfaces. For example, one can define the expected Pareto front surface as the polar surface constructed using the expected projected lengths (Example 4.1).
@@
As shown by Tu et al. (2024, Section 4.1), if the objective function $f$ is bounded and the Pareto front surfaces $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\{f(\mathbf{x},\boldsymbol{\xi})\}_{\mathbf{x}\in\mathbb{X}}]$ are non-empty almost surely, then this polar surface is a valid Pareto front surface. Otherwise it could degenerate to the singleton set $\{\boldsymbol{\eta}\}\in\mathbb{L}_{\boldsymbol{\eta}}$.
We now present a straightforward extension of this result below, which holds for any partially coherent univariate risk functional. The proof of this result is presented in Section A.6.
@@
for some positive directions $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$, where $P>0$ denotes a cardinality constraint. Individually, each of these problems tries to identify a collection of inputs $X\subseteq\mathbb{X}$, whose collective output, under the preference direction $\boldsymbol{\lambda}\in\mathcal{S}^{M-1}_{+}$, is of high quality. This problem reduces down to the STR length scalarised problem (12) when $P=1$. Notably, this is a monotone and submodular set optimisation problem when $P>1$ and therefore is, in general, much more computationally challenging to solve than the corresponding RTS or STR problems—see Tu et al. (2025) for more details.
@@ 4.3 Robust Pareto front surfaces
In Proposition 4.2 and Proposition 4.3, we present the main results in this section which states that these polar surfaces (22) and (23) are indeed valid Pareto front surfaces. Also, we give an explicit representation on what these Pareto front surfaces actually are. In words, the RTS front is the Pareto front surface associated with the union robust problem (10), whilst the STR front is the Pareto front surface associated with the risk statistic of each input’s Pareto front surface distribution. The proof of these result are presented in Section A.7 and Section A.8, respectively.
@@
In Figure 5, we illustrate these risk-adjusted Pareto front surfaces for a simple two-dimensional example with the risk functionals introduced in Section 2.3. We also compare these Pareto front surfaces with the risk statistics described in Proposition 4.1. Overall, we see a great variety between the Pareto front surfaces illustrated within each plot; the differences among the plots however are very subtle. Conceptually, the Pareto front surface statistics looks at the aggregated performance at each scalarised value and therefore has a noticeably smoother surface. In contrast, the RTS and STR front looks at the individual performance at each scalarised level and therefore is a bit more bumpy. In particular, the RTS front treats every input $\mathbf{x}\in X$ according to the Pareto front surface of its risk-adjusted output set $\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[\boldsymbol{\rho}[f(\mathbf{x},\cdot)]]$, whilst the STR front treats every input $\mathbf{x}\in X$ according to a risk statistic of its random Pareto front surface $\rho[\mathcal{Y}_{\boldsymbol{\eta}}^{\textnormal{int}}[f(\mathbf{x},\cdot)]]$. These two formulations are different in general. However, as highlighted in Section 3.3.3, there are special cases when they are equivalent, such as: the worst-case setting (blue lines) and the VaR setting (orange lines).
Figure 5: An illustration of the different Pareto front surfaces that can be computed from an $M=2$ dimensional objective function. On the left, we plot the output sets $f(\mathbf{x},\Xi)\subset\mathbb{R}^{M}$ for each input $\mathbf{x}\in\mathbb{X}$, which are all assumed to be uniformly distributed in some ellipsoid. On the remaining plots, we draw the Pareto front surfaces associated with the Pareto front statistics, RTS fronts and STR fronts for the different risk functionals described in Section 2.3.
@@ 5 Robust performance metrics
Performance metrics are often used in multi-objective optimisation in order to quantify the quality of a Pareto front approximation. As showcased in a recent survey by Audet et al. (2021), there exists many different types of performance metrics in the multi-objective literature, all of which try to measure different facets of the Pareto front approximation. Very little work has focussed on extending these existing performance metrics to the robust setting (Mirjalili and Lewis, 2015). More specifically, in this latter setting, one wishes to quantify the quality of a set of proposed inputs $X\subseteq\mathbb{X}$, based on the corresponding set of output sets $\{f(\mathbf{x},\Xi)\}_{\mathbf{x}\in X}$. Naturally, if the individual output sets are singletons, $|f(\mathbf{x},\Xi)|=1$, then we reduce down to the standard multi-objective set-up. In this section, we will now address this gap in the literature, and present two family of robust multi-objective performance metrics based on the RTS and STR approaches. Notably, we generalise the R2 utilities (Hansen and Jaszkiewicz, 1998), which are a general family of multi-objective performance metrics that exhibit many favourable properties and contains many of the most widely-used performance metrics as special cases (Tu et al., 2025). As explicit examples, we will proceed to show how one can generalise the well-known inverted generational distance (Section 5.2) and hypervolume indicator (Section 5.3) to the robust setting.
@@ 5.1 Robust R2 utilities
Formally, a multi-objective performance metric is any utility function $U:2^{\mathbb{R}^{M}}\rightarrow\mathbb{R}$, which maps a set of vectors to a scalar. We adopt the convention that larger utility values signify better performance. The R2 utilities are a special class of utility functions that is defined with respect to a family of scalarisation functions $\{s_{\boldsymbol{\theta}}:\mathbb{R}^{M}\rightarrow\mathbb{R}\}_{\boldsymbol{\theta}\in\Theta}$ and a probability distribution over the corresponding scalarisation parameters $p\in\mathcal{P}[\Theta]$. Specifically, equipped with this set-up, the corresponding R2 utility of a set of vectors $Y\subset\mathbb{R}^{M}$ is equal to the weighted average maximum scalarised value
@@
Note that this is the general form of an R2 utility which also accounts for the setting where the Pareto front approximation $Y\subset\mathbb{R}^{M}$ is continuous. Nevertheless, in practice, the Pareto front approximation is typically finite and therefore the supremum operation is often replaced with the maximum. At an intuitive level, an R2 utility just computes a weighted score associated with a collection of scalarised problems (3). By design, this family of utility functions is very flexible and it gives us any easy way to define and measure different notions of performance in the objective space. In theory, one could take advantage of this flexibility in order to define R2 utilities which adequately cater for the decision maker’s internal preferences.
@@
Inspired by the construction of the R2 utilities (24), we now define the corresponding RTS and STR extension of the R2 utilities by replacing the maximum scalarised values arising in (3) with the RTS scalarised values (8) and STR scalarised values (11), respectively. More precisely, we define an RTS-R2 utility function $R_{f,\boldsymbol{\rho}}^{\text{RTS}}:2^{\mathbb{X}}\rightarrow\mathbb{R}$ and a STR-R2 utility function $R_{f,\rho}^{\text{STR}}:2^{\mathbb{X}}\rightarrow\mathbb{R}$ by the following equations:
@@
Notice that one can rewrite an RTS-R2 utility in terms of a standard R2 utility: $R^{\textnormal{RTS}}_{f,\boldsymbol{\rho}}[X]=U[\cup_{\mathbf{x}\in X}\boldsymbol{\rho}[f(\mathbf{x},\cdot)]]$. Therefore, to compute the RTS performance of a set of inputs, one just has to compute the corresponding R2 utility of the union risk-adjusted output set. In general, a similar result does not hold for the STR-R2 utility unless the operations commute (Section 3.3.3).
@@ 5.2 Robust inverted generational distance
The inverted generational distance (IGD) is a popular multi-objective performance metric that was initially introduced by Coello and Reyes Sierra (2004). Following Schutze et al. (2012), the IGD indicator $I^{\text{IGD}_{p,q}}:2^{\mathbb{R}^{M}}\times B[\mathbb{R}^{M}]\rightarrow\mathbb{R}$ computes some notion of the average distance between a set of vectors $Y\subset\mathbb{R}^{M}$ and a finite set of ideal (or utopia) vectors $\Upsilon\in\mathbb{B}[\mathbb{R}^{M}]$,
@@
for some norms $p,q\geq 1$, where $\mathbb{B}[\mathbb{R}^{M}]:=\{Y\subseteq\mathbb{R}^{M}:|Y|<\infty\}$ denotes the space of finite sets of vectors. By construction, a smaller IGD indicator value is more desirable. Following Tu et al. (2025), we can rewrite the IGD indicator as a transformation of an R2 utility, namely the IGD utility
@@
which holds for any set of inputs $X\subseteq\mathbb{X}$ and norms $p,q\geq 1$. In other words, the RTS-IGD utility computes the average negative distance between the risk-adjusted output set and the set of utopia points, whilst the STR-IGD utility computes the average risk-adjusted negative distance between the distribution of possible objective vectors and the set of utopia points. On the left of Figure 6, we give a pictorial illustration of these different computations on a simple two-dimensional example.
@@
We can also generalise the above discussion for the corresponding IGD+ indicator described by Ishibuchi et al. (2015). This is a straightforward extension, where the IGD scalarisation function is replaced with its truncated version (Tu et al., 2025).
@@ 5.3 Robust hypervolume indicator
for any set of inputs $X\subseteq\mathbb{X}$, bounded objective function $f:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$, partially coherent univariate risk functional $\rho$ and reference vector $\boldsymbol{\eta}\in\mathbb{R}^{M}$. By a simple calculation, we see that the variance of this estimate can be uniformly upper bounded by a function dependent constant divided by the number of Monte Carlo samples, that is
@@ 6 Numerical examples
We now present two numerical case studies in order to highlight the efficacy of the novel ideas that we have presented in this work. Firstly, in Section 6.1, we present the cake baking case study which illustrates a practical example where one has to decide between an RTS and STR solution. Secondly, in Section 6.2, we present the robust rocket injector case study which showcases how one can use the concepts of a robust Pareto front and a robust performance metric in order to devise and assess the performance of some robust multi-objective optimisation algorithms.
@@ 6.1 Cake problem
changes when the relative composition of its $D=6$ main ingredients are varied. The inputs for this problem are the relative amount of ingredients $\mathbf{x}\in\mathbb{X}$, which lie in the $(D-1)$-dimensional simplex and satisfy the following inequality constraints:
@@
Suppose for instance that the decision maker is interested in identifying the ingredients that are required to make a cake whose properties achieve a target value of $\boldsymbol{\upsilon}=(-360,4,1)\in\mathbb{R}^{M}$. As motivated earlier in Example 3.1, one can recast this target optimisation problem as a robust scalarised optimisation problem. Specifically, one can use the expectation functionals (Examples 2.3 and 2.8) along with a distance-based scalarisation function. As a concrete example, we consider using the following weighted variant of the IGD scalarisation function:
@@ 6.2 Robust rocket injector problem
For the manufacturer’s distribution, we assume a uniform distribution $p(\boldsymbol{\xi})\sim\text{Uniform}(\Xi)$ over the uncertain parameters $\boldsymbol{\xi}\in\Xi$. For our robust objective, we consider an STR approach where the goal is to identify the STR robust designs whose output distribution lies on the CVaR Pareto front surface at level $\alpha=0.9$. Equivalently, we formulate this problem as a robust utility optimisation problem where we are interested in optimising the STR hypervolume (Proposition 5.2) of the identified CVaR front
@@
To emulate a real-world scenario, we treat the objective function $f$ as a black-box function that can only be evaluated sparingly at a few carefully selected design locations $T=65$. More precisely, at any time $t>0$, the decision maker is given the opportunity to see a noisy function evaluation at any chosen design $\mathbf{x}_{t}\in\mathbb{X}$, that is $\mathbf{y}_{t}=f(\mathbf{x}_{t},\boldsymbol{\xi}_{t})+\boldsymbol{\epsilon}_{t}\in\mathbb{R}^{M}$, where $\boldsymbol{\xi}_{t}\sim p(\boldsymbol{\xi})$ is the uncertain parameter that was independently sampled by the manufacturer and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{M},\textsc{Diag}(\boldsymbol{\sigma}^{2}))$ is the observation noise. In our example, we consider a zero-mean Gaussian observation noise whose standard deviation $\boldsymbol{\sigma}=(0.01,0.01,0.01)$ is set to $1\%$ of the estimated scaled objective range. To initialise this problem, we randomly sampled $T_{0}=5$ initial points from the controllable input space. In Figure 9, we illustrate these initial points in the objective space and plot the corresponding CVaR front under the STR approach. Notably, to compute this CVaR front, we require knowledge of the true objective function at each uncertain parameter, which is typically not available in a real-world set-up.
@@
where $X_{t-1}=\{\mathbf{x}_{n}\}_{n\leq t-1}$ denotes the set of previously queried inputs, whilst $\hat{\boldsymbol{\eta}}_{t}\in\mathbb{R}^{M}$ and $\hat{f}_{t-1}:\mathbb{X}\times\Xi\rightarrow\mathbb{R}^{M}$ denotes the latest estimate of the reference vector and objective function, respectively. In our example, we set the latest reference vector to be an over-estimation of the lower bound $\eta_{t-1}^{(m)}=\min_{n<t}y_{n}^{(m)}-\kappa(\max_{n<t}y_{n}^{(m)}-\min_{n<t}y_{n}^{(m)})$ with $\kappa=0.1$ for $m=1,\dots,M$. For the surrogate model, we used an upper confidence bound of a Gaussian process model which was trained on the previous observations, that is $\hat{f}^{\text{UCB}}_{t-1}(\mathbf{x},\boldsymbol{\xi})=\boldsymbol{\mu}_{t-1}(\mathbf{x},\boldsymbol{\xi})+\beta\boldsymbol{\sigma}_{t-1}(\mathbf{x},\boldsymbol{\xi})$, where $\beta=2.0$ is the trade-off parameter and the vectors $\boldsymbol{\mu}_{t-1}(\mathbf{x},\boldsymbol{\xi})\in\mathbb{R}^{M}$ and $\boldsymbol{\sigma}_{t-1}(\mathbf{x},\boldsymbol{\xi})\in\mathbb{R}_{\geq 0}^{M}$ are the mean and standard deviation of the model at the input pair $(\mathbf{x},\boldsymbol{\xi})\in\mathbb{X}\times\Xi$. We refer to this overall acquisition strategy as the STR-HVI-UCB approach.
@@
Figure 10: A numerical comparison between the different multi-objective algorithms on the robust rocket injector problem over $100$ independent runs. For each method, we plot the mean log regret at every iteration along with an uncertainty estimate. Specifically, we highlight the band between the $0.25$ to $0.75$ and $0.1$ to $0.9$ quantiles with a dark and light shade, respectively.
In Figure 10, we present the results of our numerical comparison in which we compared our proposed method versus the existing baselines over $100$ independent runs on a discretised version of the robust rocket injector problem. Namely, we compare our method against non-robust [^5] These approaches are non-robust in the sense that they can condition on the observed data set $\mathcal{D}_{t-1}=\{(\mathbf{x}_{n},\boldsymbol{\xi}_{n},\mathbf{y}_{n})\}_{n\leq t-1}$, at every time $t$, but they do not consider the uncertainty present in the uncontrollable variable $\boldsymbol{\xi}\in\Xi$. approaches such as: the random search (Sobol), NSGA2 (Deb et al., 2002), NParEGO (Knowles, 2006; Daulton et al., 2021), NEHVI (Emmerich et al., 2006; Daulton et al., 2021), PFES (Suzuki et al., 2020) and the JES-LB (Tu et al., 2022) algorithms. We also compare our approach against the MARS algorithm (Daulton et al., 2022), which is a robust algorithm that takes an STR approach using the Chebyshev scalarisation function and the univariate VaR risk functional. Unsurprisingly we see that the robust methods perform better than the non-robust methods. Moreover, our method performs the best overall under the chosen utility. For some visual intuition, we have included Figure 11 where we plot the final CVaR front obtained by each algorithm under the same random seed. Visually we see that most of the variation between these algorithms transpire around the first and second objectives. The existing approaches regularly struggle to find the inputs which robustly maximises the CVaR front over these two objectives. In contrast, our robust approach is much more consistent and is often very quick in identifying the difficult inputs which robustly maximise the whole CVaR front.